# New Insights Into Achievable Spectral Efficiency With Adaptive Free Space Optical Transmissions

### v8

@@ III-A Exact ASE With and Without Pointing Errors
In the limiting case where $A_{0}\to 1$ and $\xi^{2}\to\infty$, the fading distribution converges to the pure $G$ distribution. Substituting these limits into (1) and applying the identity from [36, Eq. (9.31.1)] completes the proof of the theorem. ∎
@@ III-B Immediate Insights
  • At high SNRs, contribution from the Meijer-G based term vanishes while $\mu_{k}\to 1/\mathrm{SNR}_{k}$. Keeping this in mind, and now comparing the ASE of channels with and without pointing error, the ASE loss due to pointing error can be easily quantified in terms of $A_{0}$ and $\xi^{2}$ parameters.
@@ III-C Low-SNR and High-SNR Asymptotic Expressions
The asymptotic expansions of the exact ASE expression reveal how key channel parameters from GG turbulence and pointing error influence the performance of adaptive FSO schemes. To this end, the $\mu_{k}$$\mathrm{SNR}_{k}$ relationship is first derived under average power constraint in both low-SNR and high-SNR regimes.
@@
The threshold value $\mu_{k}$ associated with the optimal beam power adaptation as described in (16) for the transmitted laser beam satisfies
  • 2.
    At  low SNR: $\mu_{k}\,\approx\,\dfrac{A_{0}}{4ab}\ln^{2}\left(\dfrac{1}{\mathrm{SNR}_{k}}\right)$.
@@
    Part 1): Substituting (7) into the average optical power constraint (17), we have
    $\dfrac{\mu_{k}\mathrm{SNR}_{k}}{A}=\int_{\mu_{k}}^{\infty}\left(1-\dfrac{\mu_{k}}{\lambda}\right)G_{1,3}^{3,0}\left(\frac{ab\lambda}{A_{0}}\biggr{\rvert}\hskip-3.50006pt\begin{array}[]{c}\xi^{2}\\[1.00006pt] \xi^{2}-1,a-1,b-1\\ \end{array}\hskip-5.0pt\right)d\lambda,$
    where $A\,=\,{ab\xi^{2}/(A_{0}\Gamma(a)\Gamma(b)})$. Applying the identity in (22) to the power constraint equation gives
    $\dfrac{\mu_{k}\mathrm{SNR}_{k}}{A}\,=\,\dfrac{1}{A}\,$-\, $\dfrac{A_{0}}{ab}G_{2,4}^{3,1}\left(\frac{ab}{A_{0}}\mu_{k}\,\biggr{\rvert}\hskip-1.49994pt\begin{array}[]{c}1,\,\xi^{2}+1\\[1.00006pt] \xi^{2},\,a,\,b,\,0\\ \end{array}\hskip-5.0pt\right)$ (64)
    $+\,\,\mu_{k}\,$ $G_{2,4}^{3,1}\left(\dfrac{ab}{A_{0}}\mu_{k}\,\biggr{\rvert}\hskip-1.49994pt\begin{array}[]{c}1,\xi^{2}\\[1.00006pt] \xi^{2}-1,a-1,b-1,0\end{array}\hskip-5.0pt\right).$ (67)
    From the power constraint in (17), it follows that $\mu_{k}$ exhibits a strictly monotonic inverse relation with $\mathrm{SNR}_{k}$, decreasing as $\mathrm{SNR}_{k}$ increases. Notably, at high SNRs, the contributions of the Meijer-G function terms in (64) become negligible due to their vanishing behavior for small arguments. Hence
    $\lim_{\mu_{k}\to 0}\,\,\mu_{k}\mathrm{SNR}_{k}\,\,=\,\,1.$
    This completes the proof of 1)
    Part 2): We now derive an explicit expression for the $\mu_{k}$$\mathrm{SNR}_{k}$ dependence in the low-SNR regime. As $\mathrm{SNR}_{k}$ decreases, the cutoff $\mu_{k}$ increases, leading the transmitter to allocate power only for high channel gains. To facilitate this, the channel gain distribution $f_{\lambda}(\lambda)$ in (7) is approximated using a low-order series expansion of the Meijer $G$-function for large arguments, as detailed below:
    $G_{1,3}^{3,0}\left(z\,\biggr{\rvert}\hskip-1.49994pt\begin{array}[]{c}\xi^{2}\\ \xi^{2}-1,\,a-1,\,b-1\\ \end{array}\hskip-5.0pt\right)\approx e^{-2\sqrt{z}}\sqrt{\pi}z^{\frac{1}{4}(-7+2a+2b)}.$ (70)
@@
    $\mathrm{SNR}_{k}\,$ $\approx\,K\left(\dfrac{B_{1}(ab\mu_{k}/A_{0})}{ab\mu_{k}/A_{0}}-B_{2}(ab\mu_{k}/A_{0})\right)$ (71)
    where $K=(\sqrt{\pi}\,ab\,\xi^{2})/(A_{0}\,\Gamma(a)\,\Gamma(b))$ is a positive constant independent of the threshold value, and
@@
    $\mathrm{SNR}_{k}\,\approx\,K\left(\dfrac{ab}{A_{0}}\mu_{k}\right)^{\left(\frac{a+b}{2}\ -\frac{11}{4}\right)}e^{-2\sqrt{\frac{ab}{A_{0}}\mu_{k}}}.$ (75)
@@
    $\ln(\mathrm{SNR}_{k})\,\approx\,-2\sqrt{{ab\mu_{k}}/{A_{0}}}.$ (76)
    Solving (76) for $\mu_{k}$, we finally arrive at
    $\mu_{k}\,\approx\,\dfrac{A_{0}}{4ab}\ln^{2}\left(\frac{1}{\mathrm{SNR}_{k}}\right).\vskip-10.00002pt$ (77)
@@
    For the GG turbulence channel with pointing error, the asymptotic ASE at low and high SNR are given by
@@
    where the subscripts ‘lowlow’ and ‘highhigh’ stand for the low-SNR (asymptotically 0) and high-SNR (asymptotically $\infty$) regimes respectively.
@@
    Part 1): At high SNR, the threshold behaves as $\mu_{k}\approx\mathrm{SNR}_{k}^{-1}$ (from Lemma 1), while the Meijer’s G function becomes negligible for small input arguments, asymptotically approaching zero as
@@
    Part 2): The low-order series expansion of the Meijer-G function in (1) for large input argument $z$ (asymptotically $\infty$) is
    $G_{3,5}^{3,2}\left(z\ \biggr{\rvert}\ \begin{matrix}1,1,\xi^{2}+1\\ \xi^{2},a,b,0,0\end{matrix}\right)\,\approx\,e^{-2\sqrt{z}}\sqrt{\pi}\,z^{\frac{1}{4}(-7+2a+2b)}\\[3.50006pt] +\dfrac{\Gamma(a)\Gamma(b)}{\xi^{2}}\left(\ln(z)+\dfrac{1}{\xi^{2}}-\psi(a)-\psi(b)\right).$ (83)
@@
    Finally, substituting $\mu_{k}$ from  (77) in (85) completes the proof of the low SNR ASE expansion (78) in Theorem 2. ∎
@@
    • By applying $A_{0}\to 1$ and $\xi^{2}\to\infty$, the asymptotic ASEs of the GG turbulence channel ‘without pointing error’ at low and high SNR are computed as
@@
      $\dfrac{1}{k}\biggl{[}\,\dfrac{1}{\xi^{2}}\,-\,\ln A_{0}\,\biggr{]}.$ (88)
    • For the GG turbulence channel, the presence of pointing errors decreases the ASE at low SNRs by $A_{0}$ times.
@@
    In this section, we present the analytical utility of the derived ASE formulas, supported by extensive numerical results. Previous studies on related topics [25], [28], [29], [31], and [32] have suggested, both directly and indirectly, that a wide range of optical turbulence conditions can be achieved by varying the FSO link length. However, link length variations impact the large-scale path loss factor. Additionally, the laser beam waist expands with propagation distance, impacting the pointing error distribution parameters $\xi$ and $A_{0}$. Notably, these studies did not consider these interdependencies. Understanding the interactions among these channel characteristics is essential, as this work aims to accurately analyze the impact of variations in turbulence and pointing error on the ASE performance of terrestrial FSO systems.
@@
     Ch. Setting (a): Gamma-Gamma (GG) fading ‘without’ pointing errorxxxx                Turbulence   Rytov       GG strength   variance $\sigma_{\textrm{R}}^{2}$         distribution parameters Weak   $0.2$    a=  11.651,b= 10.122a\,=\,\,11.651,\,\,b\,=\,10.122 Moderate   11    a=  4.3939,b= 2.5636a\,=\,\,4.3939,\,\,b\,=\,2.5636 Strong   44    a=  4.3407,b= 1.3088a\,=\,\,4.3407,\,\,b\,=\,1.3088  Ch. Setting (b): GG fading with ‘mild’ pointing error ($\sigma_{\mathrm{e}}=1$ cm)x                Turbulence    Rytov        Pointing Error strength    variance $\sigma_{\textrm{R}}^{2}$       distribution parameters Weak $0.2$   $\xi\,=\,1.8102,\,A_{0}\,=\,0.1526$ Moderate 11   $\xi\,=\,2.1977,\,A_{0}\,=\,0.1035$ Strong 44   $\xi\,=\,3.5510,\,A_{0}\,=\,0.0397$  Ch. Setting (c): GG fading with ‘strong’ pointing error ($\sigma_{\mathrm{e}}=4$ cm)x                Turbulence    Rytov        Pointing Error strength    variance $\sigma_{\textrm{R}}^{2}$       distribution parameters Weak $0.2$   $\xi\,=\,0.4525,\,A_{0}\,=\,0.1526$ Moderate 11   $\xi\,=\,0.5494,\,A_{0}\,=\,0.1035$ Strong 44   $\xi\,=\,0.8877,\,A_{0}\,=\,0.0397$ To ensure a fair comparison, the FSO Tx-Rx system parameters (e.g., optical wavelength, Tx. beam waist, Rx. aperture) are kept fixed. In addition, a horizontal terrestrial FSO link of ‘fixed length’ is considered, ensuring that the path loss factor remains unchanged. A horizontal distance of 300 m, representative of short-to-medium haul FSO links, is selected to capture realistic turbulence and pointing error effects while keeping other impairments negligible (see Remark 1). To explicitly account for the acquisition, tracking, and pointing (ATP) subsystem, three representative GG channel settings are considered: (a) No pointing error, corresponding to an ideal or “genie-aided” ATP subsystem that fully eliminates jitter, serving as a turbulence-limited benchmark; (b) Mild pointing error, corresponding to a realistic, well-designed ATP subsystem with small residual jitter; and (c) Strong pointing error, corresponding to a poorly designed or absent ATP subsystem (i.e., no ATP), where jitter dominates. Although such severe jitter exceeds typical values observed in practical systems with ATP, it is included for theoretical completeness, highlighting the transition from turbulence-limited to jitter-dominated operation.
@@
    $\sigma_{\mathrm{R}}^{2}=1.23C_{\mathrm{n}}^{2}k_{\mathrm{w}}^{7/6}{L^{11/6}}$ (123)
@@
    For near-ground horizontal propagation, the parameter $C_{\mathrm{n}}^{2}$ remains relatively constant and primarily varies with altitude-being stronger near the ground and weaker at higher elevations [11, Sec. 12.2][50, Ch. 2].
@@
    • Recall from (16) and (17) that $\mathrm{SNR}:=P_{\mathrm{avg}}/N_{0}$ is the average transmit SNR in all the derived ASE formulae (with subscript $k$ now suppressed from the SNR symbol as the noise levels are assumed the same) and that the horizontal axis in Figures 1-7 is the transmit SNR in log scale, not the received SNR. This enables fair comparison and accurate analysis mentioned earlier. If required, the average received SNR can be computed by scaling the transmit $\mathrm{SNR}$ by the mean channel gain $\mathbb{E}[\lambda]$.
@@
    • As discussed in the introduction, a fundamental distinction between the two FSO wireless systems, namely IM-DD and coherent HD-based communication systems, lies in their noise characteristics [10]. For the same transmit SNR, the received SNR after photodetection is typically higher with a coherent receiver than with an IM-DD receiver. For analytical consistency and concreteness, we assume that the noise power levels (per Hz) in both detection receivers are the same, $N_{1}=N_{2}=N_{0}$ say.
    1) Exact ASE behavior: Based on Theorem 1, the exact ASE for HD and IM-DD optical channels under GG turbulence, with and without pointing errors, is computed and shown in Figures 1 and 2, respectively. The analytical results closely match Monte Carlo simulation outcomes across a wide SNR range and under various turbulence strengths and pointing jitter conditions.
    Figure 1: Average spectral efficiency of the GG turbulence channel without pointing error under HD and IM-DD detection schemes. The average received SNR is the same as the transmit SNR since the mean channel gain $\mathbb{E}[\lambda]\,=\,1$.
    Figure 2: Average spectral efficiency of the GG turbulence channel with pointing jitter under coherent HD detection scheme. The received SNR is less than the transmit SNR by the mean channel gain $\mathbb{E}[\lambda]$.
@@
    • As shown in Fig. 1, spectral efficiency deteriorates with increasing turbulence at high SNR for both detection schemes. This is expected, as stronger turbulence leads to a heavier concentration of low channel gains in the PDF (cf. Fig. 6).
@@
    2) ASE behavior at High SNRs: The high-SNR asymptotic ASE behavior in (79), as established in Theorem 2, invites comparison with the ASE of an ideal AWGN channel, which is asymptotically given by $\ln\mathrm{SNR}$. The resulting difference-being strictly negative-quantifies the spectral efficiency loss at high SNRs due to the combined impairments of GG turbulence and pointing jitter, i.e.,
[…]
    Figure 3: The loss in average spectral efficiency at high SNRs as a function of turbulence strength for the GG channel with pointing jitter under HD detection scheme.
    Discussion: Fig. 3 illustrates the spectral efficiency penalty (124) plotted as a function of atmospheric turbulence strength for mild and strong pointer error channel settings (see Table IV). Recall that a fixed-length horizontal FSO link is considered, with a laser beam of waist $w_{0}=4.5$ mm at the transmitter (see Table IV). The beam footprint at the receiver is broadened due to beam divergence caused by turbulence over the link length. For the strong pointing error channel setting, the enlarged beam footprint leads to
    • a reduction in collected power at the fixed-aperture receiver (i.e., decreased $A_{0}$); and
@@
    As shown in Fig. 3, for the GG fading with strong pointing error channel setting, the ASE at high SNR increases with turbulence in the practical $0<\sigma_{\mathrm{R}}^{2}\leq 4$ range. This ASE behavior arises because the reduction in the pointing jitter effect more than compensates for the loss in collected power at the receiver. Consequently, the ASE penalty spans approximately 10-10 to 7.1-7.1 bits/s/Hz, indicating a remarkable improvement of nearly $3.0$ bits/s/Hz in the high SNR regime.
    For the GG fading with mild pointing error channel setting, no such trade-off is observed, as $\xi$ exhibits negligible turbulence-induced gain while $A_{0}$ continues to decay, leading to a monotonic ASE loss. The corresponding high-SNR penalty, illustrated by the blue curve in Fig. 3, ranges from 3.2-3.2 to 6.1-6.1 bits/s/Hz.
@@
[…]
    which corresponds to a relatively small reduction, as can be inferred from Fig. 2 (further qualification in Fig. 4 below).
    Figure 4: Average spectral efficiency of the GG turbulence channel ”without” pointing error under HD scheme: $A_{0}=1$ for “ideal” system condition whereas $A_{0}$ from Table IV are considered as representation for realistic system condition.

### v9

@@ III-A Exact ASE With and Without Pointing Errors
In the limiting case where $A_{\mathrm{\scriptscriptstyle 0}}\to 1$ and $\xi^{2}\to\infty$, the fading distribution converges to the pure $G$ distribution. Substituting these limits into (1) and applying the identity from [36, Eq. (9.31.1)] completes the proof of the theorem. ∎
@@ III-B Immediate Insights
  • At high SNRs, contribution from the Meijer-G based term vanishes while $\mu\to 1/\mathrm{SNR}$. Keeping this in mind, and now comparing the ASE of channels with and without pointing error, the ASE loss due to pointing error can be easily quantified in terms of $A_{\mathrm{\scriptscriptstyle 0}}$ and $\xi^{2}$ parameters.
@@ III-C Low-SNR and High-SNR Asymptotic Expressions
[…]
@@
The threshold value $\mu$ associated with the optimal beam power adaptation as described in (14) for the transmitted laser beam satisfies
$\mathrm{(i)}\,\,\,\text{At high SNR:}\,\,\mu\,\approx\,\dfrac{1}{\mathrm{SNR}},$ (72)
$\mathrm{(ii)}\,\,\text{At low SNR:}\,\,\,\mu\,\approx\,\dfrac{A_{\mathrm{\scriptscriptstyle 0}}}{4ab}\ln^{2}\left(\dfrac{1}{\mathrm{SNR}}\right).$ (73)
@@
Part (i): Substituting (9) into the average optical power constraint (15), we have
$\dfrac{\mu\mathrm{SNR}}{A}=\int_{\mu}^{\infty}\biggl(1-\dfrac{\mu}{\lambda}\biggr)\,G_{1,3}^{3,0}\left(\frac{ab\lambda}{A_{\mathrm{\scriptscriptstyle 0}}}\biggr\rvert\hskip-3.50006pt\begin{array}[]{c}\xi^{2}\\[1.00006pt] \xi^{2}-1,a-1,b-1\\ \end{array}\hskip-3.50006pt\right)d\lambda,$
where $A\,:=\,{ab\xi^{2}/(A_{\mathrm{\scriptscriptstyle 0}}\Gamma(a)\Gamma(b)})$. Applying the identity in (20) to the power constraint equation gives
$\dfrac{\mu\mathrm{SNR}}{A}\,=\,\dfrac{1}{A}\,$-\, $\dfrac{A_{\mathrm{\scriptscriptstyle 0}}}{ab}G_{2,4}^{3,1}\left(\frac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\,\xi^{2}+1\\[1.00006pt] \xi^{2},\,a,\,b,\,0\\ \end{array}\hskip-3.50006pt\right)$ (76)
$+\,\,\mu\,$ $G_{2,4}^{3,1}\left(\dfrac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\xi^{2}\\[1.00006pt] \xi^{2}-1,a-1,b-1,0\end{array}\hskip-3.50006pt\right).$ (79)
From the power constraint in (15), it follows that $\mu$ exhibits a strictly monotonic inverse relation with $\mathrm{SNR}$, decreasing as $\mathrm{SNR}$ increases. Notably, at high SNRs, the contributions of the Meijer-G function terms in (76) become negligible due to their vanishing behavior for small arguments. Hence limμ→0    μSNR   =   1. This completes the proof of 1).
[…]
$G_{1,3}^{3,0}\left(z\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}\xi^{2}\\ \xi^{2}-1,a-1,b-1\\ \end{array}\hskip-3.50006pt\right)\,\approx\,\dfrac{\sqrt{\pi}\,z^{(-7+2a+2b)/4}}{e^{2\sqrt{z}}}.$ (82)
@@
$\mathrm{SNR}\,$ $\approx\,\mathcal{K}\left(\dfrac{B_{1}(ab\mu/A_{\mathrm{\scriptscriptstyle 0}})}{ab\mu/A_{\mathrm{\scriptscriptstyle 0}}}-B_{2}(ab\mu/A_{\mathrm{\scriptscriptstyle 0}})\right)$ (83)
where $\mathcal{K}:=(\sqrt{\pi}\,ab\,\xi^{2})/(A_{\mathrm{\scriptscriptstyle 0}}\,\Gamma(a)\,\Gamma(b))$ is a positive constant independent of the threshold value, and
@@
$\mathrm{SNR}\,\approx\,\mathcal{K}\left(\dfrac{ab\mu}{A_{\mathrm{\scriptscriptstyle 0}}}\right)^{\left(\frac{a+b}{2}-\frac{11}{4}\right)}e^{-2\sqrt{\frac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu}}.$ (87)
@@
$\ln(\mathrm{SNR})\,\approx\,-2\sqrt{{ab\mu}/{A_{\mathrm{\scriptscriptstyle 0}}}}.$ (88)
Solving (88) for $\mu$, we finally arrive at
$\mu\,\approx\,\dfrac{A_{\mathrm{\scriptscriptstyle 0}}}{4ab}\ln^{2}\left(\frac{1}{\mathrm{SNR}}\right).\vskip-10.00002pt$ (89)
@@
For the coherent HD optical channel subjected to GG turbulence with pointing errors, the asymptotic ASE at low and high SNR are given by
@@
where the superscripts $\mathrm{low}$’ and $\mathrm{high}$’ stand for the low-SNR (asymptotically 0) and high-SNR (asymptotically $\infty$) regimes respectively.
@@
Part (i): At high SNR, the threshold behaves as $\mu\approx\mathrm{SNR}^{-1}$ (from Lemma 1), while the Meijer’s G function becomes negligible for small input arguments, asymptotically approaching zero as
@@
Part (ii): The low-order series expansion of the Meijer-G function in (1) for large input argument $z$ (asymp. $\infty$) is
$G_{3,5}^{3,2}\left(z\,\biggr\rvert\,\,\begin{matrix}1,1,\xi^{2}+1\\ \!\xi^{2},a,b,0,0\end{matrix}\right)\approx\,$ $\,\dfrac{\sqrt{\pi}\,e^{-2\sqrt{z}}}{z^{(7-2a-2b)/4}}\,+\,\dfrac{\Gamma(a)\,\Gamma(b)}{\xi^{2}}\,\times$ (93)
$\left[\ln(z)-\psi(a)-\psi(b)+\dfrac{1}{\xi^{2}}\right].$ (94)
@@
Finally, substituting $\mu$ from  (89) in (96) completes the proof of the low SNR ASE expansion (90) in Theorem 2. ∎
@@
  • By applying $A_{\mathrm{\scriptscriptstyle 0}}\to 1$ and $\xi^{2}\to\infty$, the asymptotic ASEs of the GG turbulence channel ‘without pointing errors’ at low and high SNR are computed as
@@
    $\biggl[\dfrac{1}{\xi^{2}}\,-\,\ln A_{\mathrm{\scriptscriptstyle 0}}\biggr].$ (99)
  • For the GG turbulence channel, the presence of pointing errors decreases the ASE at low SNRs by $A_{\mathrm{\scriptscriptstyle 0}}$ times.
@@ IV Numerical Results and Discussion
In this section, we present the analytical utility of the derived ASE formulas, supported by extensive numerical results. Previous studies on related topics [25], [28], [29], [31], and [32] have suggested, both directly and indirectly, that a wide range of optical turbulence conditions can be achieved by varying the FSO link length. However, link length variations impact the large-scale path loss factor. Additionally, the laser beam waist expands with propagation distance, impacting the pointing error distribution parameters $\xi$ and $A_{\mathrm{\scriptscriptstyle 0}}$. Notably, these studies did not consider these interdependencies. Understanding the interactions among these channel characteristics is essential, as this work aims to accurately analyze the impact of variations in turbulence and pointing error on the ASE performance of terrestrial FSO systems.
@@
Parameter       Symbol Value
Optical wavelength       $\lambda_{\mathrm{w}}$   1550 nm
Tx. Beam waist       $w_{\mathrm{\scriptscriptstyle 0}}$  0.45 cm
FSO channel length       $L$  300 m
Rx. Aperture radius       $r_{\mathrm{\scriptscriptstyle A}}$  1 cm
Tx./Rx. optics eff.       -  100% (assumed)
Path loss factor       -    Unity (assumed)
 Ch. Setting (a): Gamma-Gamma (GG) fading ‘without’ pointing errorxxxx
Turbulence   Rytov       $G$
strength   variance $\sigma_{{\mathrm{\scriptscriptstyle R}}}^{2}$         distribution parameters
Weak   $0.2$    a=  11.651,b= 10.122a\,=\,\,11.651,\,\,b\,=\,10.122
Moderate   11    a=  4.3939,b= 2.5636a\,=\,\,4.3939,\,\,b\,=\,2.5636
Strong   44    a=  4.3407,b= 1.3088a\,=\,\,4.3407,\,\,b\,=\,1.3088
 Ch. Setting (b): GG fading with ‘mild’ pointing error ($\sigma_{\mathrm{e}}=1$ cm)x
Turbulence    Rytov        Pointing Error
strength    variance $\sigma_{{\mathrm{\scriptscriptstyle R}}}^{2}$       distribution parameters
Weak $0.2$   $\xi\,=\,1.8102,\,A_{\mathrm{\scriptscriptstyle 0}}\,=\,0.1526$
Moderate 11   $\xi\,=\,2.1977,\,A_{\mathrm{\scriptscriptstyle 0}}\,=\,0.1035$
Strong 44   $\xi\,=\,3.5510,\,A_{\mathrm{\scriptscriptstyle 0}}\,=\,0.0397$
 Ch. Setting (c): GG fading with ‘strong’ pointing error ($\sigma_{\mathrm{e}}=4$ cm)x
Turbulence    Rytov        Pointing Error
strength    variance $\sigma_{{\mathrm{\scriptscriptstyle R}}}^{2}$       distribution parameters
Weak $0.2$   $\xi\,=\,0.4525,\,A_{\mathrm{\scriptscriptstyle 0}}\,=\,0.1526$
Moderate 11   $\xi\,=\,0.5494,\,A_{\mathrm{\scriptscriptstyle 0}}\,=\,0.1035$
Strong 44   $\xi\,=\,0.8877,\,A_{\mathrm{\scriptscriptstyle 0}}\,=\,0.0397$
To ensure a fair comparison, the FSO Tx-Rx system parameters (e.g., optical wavelength, Tx. beam waist, Rx. aperture) are kept fixed. In addition, a horizontal terrestrial FSO link of ‘fixed length’ is considered, ensuring that the path loss factor remains unchanged. A horizontal distance of 300 m, representative of short-to-medium haul FSO links, is selected to capture realistic turbulence and pointing error effects while keeping other impairments negligible (see Remark 1). To explicitly account for the acquisition, tracking, and pointing (ATP) subsystem, three representative GG channel settings are considered: (a) No pointing error, corresponding to an ideal or “genie-aided” ATP subsystem that fully eliminates jitter, serving as a turbulence-limited benchmark; (b) Mild pointing error, corresponding to a realistic, well-designed ATP subsystem with small residual jitter; and (c) Strong pointing error, corresponding to a poorly designed or absent ATP subsystem (i.e., no ATP), where jitter dominates. Although such severe jitter levels exceed typical values observed in practical FSO systems employing ATP, they are nevertheless included for theoretical completeness, as they highlight the transition from turbulence-limited to jitter-dominated operation.
@@
$\sigma_{{\mathrm{\scriptscriptstyle R}}}^{2}=1.23C_{\mathrm{n}}^{2}k_{\mathrm{w}}^{7/6}{L^{11/6}}$ (100)
@@
For near-ground horizontal propagation, the parameter $C_{\mathrm{n}}^{2}$ remains relatively constant and primarily varies with altitude-being stronger near the ground and weaker at higher elevations [11, Sec. 12.2][48, Ch. 2].
@@
  • Recall from (14) and (15) that $\mathrm{SNR}:=P_{\mathrm{avg}}/N_{\mathrm{hd}}$ is the average transmit SNR in all the derived ASE formulae, and that the horizontal axis in Figures 1-7 is the transmit SNR in log scale, not the received SNR. This enables fair comparison and accurate analysis mentioned earlier. If required, the average received SNR can be computed by scaling the transmit $\mathrm{SNR}$ by the mean channel gain $\mathbb{E}[\lambda]$.
@@
1) Exact ASE behavior: Based on Theorem 1, the exact ASE for HD-detection based coherent optical channel under GG turbulence, with and without pointing errors, is computed and shown in Figures 1 and 2, respectively. The analytical results closely match Monte Carlo simulation outcomes across a wide SNR range and under various turbulence strengths and pointing jitter conditions.
Figure 1: Average spectral efficiency of the GG turbulence channel without pointing errors under the coherent HD detection scheme. The average received SNR is the same as the transmit SNR since the mean channel gain $\mathbb{E}[\lambda]\,=\,1$.
Figure 2: Average spectral efficiency of the GG turbulence channel with pointing jitter under the coherent HD detection scheme. The received SNR is less than the transmit SNR by the mean channel gain $\mathbb{E}[\lambda]$.
@@
  • As shown in Fig. 1, spectral efficiency deteriorates with increasing turbulence at high SNR. This is expected, as stronger turbulence leads to a heavier concentration of low channel gains in the PDF (cf. Fig. 6).
@@
2) ASE behavior at High SNRs: The high-SNR asymptotic ASE behavior in (91), as established in Theorem 2, invites comparison with the ASE of an ideal AWGN channel, which is asymptotically given by $\ln\mathrm{SNR}$. The resulting difference—being strictly negative—quantifies the spectral efficiency loss at high SNRs due to the combined impairments of GG turbulence and pointing jitter, and is given by
$\Delta\overline{\mathrm{S}}_{\mathrm{\scriptscriptstyle HD}}\,\triangleq\,\ln\left(\frac{A_{\mathrm{\scriptscriptstyle 0}}}{ab}\right)\,+\,\psi(a)\,+\,\psi(b)\,-\,\left(\frac{1}{\xi^{2}}\right).$ (101)
Discussion: Fig. 3 depicts the spectral efficiency penalty (101) plotted as a function of atmospheric turbulence strength for the mild and strong pointer error channel settings in Table I. Recall that a fixed-length horizontal FSO link is considered, with a laser beam of waist $w_{\mathrm{\scriptscriptstyle 0}}=4.5$ mm at the transmitter (see Table I). The beam footprint at the receiver is broadened due to beam divergence caused by turbulence over the link length. For the strong pointing error channel setting, the enlarged beam footprint leads to
  • a reduction in collected power at the fixed-aperture receiver (i.e., decreased $A_{\mathrm{\scriptscriptstyle 0}}$); and
@@
Figure 3: The average spectral efficiency loss (in magnitude) at high SNRs as a function of turbulence strength for the GG channel with pointing jitter under the coherent HD detection scheme.
As shown in Fig. 3, for the GG fading with strong pointing error channel setting, the ASE at high SNR increases with turbulence in the practical $0<\sigma_{\mathrm{\scriptscriptstyle R}}^{2}\leq 4$ range. This atypical ASE behavior with turbulence arises because the reduction in the pointing jitter effect more than compensates for the loss in collected power at the receiver. Consequently, the ASE penalty (magnitude) decreases approximately from $10$ to $7.1$ bits/s/Hz, indicating a remarkable improvement of nearly $3.0$ bits/s/Hz in the high SNR regime.
For the GG fading with mild pointing error channel setting, no such trade-off is observed, as increased $\xi$ induces marginal improvement while $A_{\mathrm{\scriptscriptstyle 0}}$ continues to decay significantly, leading overall to a monotonic ASE loss with turbulence. The corresponding high-SNR penalty magnitude, illustrated by the blue curve in Fig. 3, ranges from $3.2$ to $6.1$ bits/s/Hz.
@@
$\Delta\overline{\mathrm{S}}_{\mathrm{\scriptscriptstyle HD}}\,=\,\psi(a)\,+\,\psi(b)\,-\,\ln\left(ab\right)$ (102)
which corresponds to a relatively small reduction, as can be inferred from Fig. 2 (further qualification in Fig. 4).
Figure 4: Average spectral efficiency of the GG turbulence channel without pointing error under the coherent HD scheme: $A_{\mathrm{\scriptscriptstyle 0}}=1$ for “ideal” system condition whereas $A_{\mathrm{\scriptscriptstyle 0}}$ from Table I are considered as representation for realistic system condition.